please help me solve this sequence of numbers
fill in the blank 2, 8,27,85,260,________,2365
i have tried adding,,,,,multiplying....i'm stuck

Answer by scott8148(6628)

try multiplying and adding ... the multiplier is constant but the addend is the term number

2*3+2=8
8*3+3=27
27*3+4=85
85*3+5=260

260*3+6=786

786*3+7=2365